# Learning-Based Modeling of a Magnetically Steerable Soft Suction Device for Endoscopic Endonasal Interventions

### v2

@@ Abstract
This paper introduces a novel learning-based modeling framework for a magnetically steerable soft suction device designed for endoscopic endonasal brain tumor resection. The device is miniaturized (4 mm outer diameter, 2 mm inner diameter, 40 mm length), 3D printed using biocompatible SIL 30 material, and integrates embedded Fiber Bragg Grating (FBG) sensors for real-time shape feedback. Shape reconstruction is represented using four Bezier control points, allowing for a compact and smooth representation of the device’s deformation. A data-driven model was trained on 5,097 experimental samples covering a range of magnetic field magnitudes (0–14 mT), actuation frequencies (0.2–1.0 Hz), and vertical tip distances from the surface of the electromagnet coil table (90–100 mm), using both Neural Network (NN) and Random Forest (RF) architectures. The RF model outperformed the NN across all metrics, achieving a mean root mean square error of 0.087 mm in control point prediction and a mean shape reconstruction error of 0.064 mm. Feature importance analysis further revealed that magnetic field components predominantly influence distal control points, while frequency and distance affect the base configuration. This learning-based approach effectively models the complex nonlinear behavior of hyperelastic soft robots under magnetic actuation without relying on simplified physical assumptions. By enabling sub-millimeter shape prediction accuracy and real-time inference, this work represents an advancement toward the intelligent control of magnetically actuated soft robotic tools in minimally invasive neurosurgery.
@@ 1 Introduction
Figure 1: Magnetically steerable soft suction device for endoscopic endonasal brain tumor resection. The device is introduced through the nasal cavity and sphenoid sinus using a conventional suction endoscope to access deep-seated regions at the skull base. A hollow cylindrical permanent magnet at the tip of the soft robot enables remote steering via externally applied magnetic fields. This magnetic actuation allows the device to navigate around critical anatomical structures, such as the pituitary gland, to precisely target and extract tumor tissue (e.g., tuberculum sellae meningioma).
@@
Recognizing the potential of combining soft robotics and magnetic actuation, a prior study in [23] demonstrated the feasibility of integrating a magnetically steerable soft tip with a conventional rigid suction device to overcome the limitations of straight, rigid tools in endoscopic procedures. The prototype in the previous study featured soft segments made from Ecoflex 00-50 embedded with a permanent magnet, and it achieved controlled in-plane bending angles exceeding 90 under moderate magnetic fields (below 25 mT). The device was fabricated using mold-based casting techniques, offering both sufficient suction capability and flexibility. Validation experiments in simulated neurosurgical environments, using agar-based brain phantoms, demonstrated that the device could safely access target sites and perform precise tumor extraction, highlighting its potential to improve the safety and efficacy of minimally invasive neurosurgical interventions.
@@
Shape estimation in flexible robots can generally be achieved using either sensor-free or sensor-based methods. Sensor-free approaches rely on kinematic or magnetic models to infer shape without embedded sensors [28, 24], while sensor-based techniques use embedded elements for direct measurement [25, 26]. Our method follows the sensor-based approach, combining FBG data with a learning-based model to enable accurate, real-time shape prediction. More importantly, a learning-based modeling framework is proposed to replace the traditional mechanical models. This data-driven approach directly maps magnetic field inputs to the device’s shape using Bezier curve control points, effectively capturing the nonlinear deformation behavior without relying on oversimplified physical assumptions on the magnetic field actuation and robotic dynamics. This enables precise shape prediction, improved control, and adaptability across diverse surgical scenarios. To the best of our knowledge, this is the first magnetically actuated soft suction device equipped with a learning-based shape prediction model tailored for neurosurgical applications. The key contributions are:
@@
    Development of a data-driven modeling framework that directly maps magnetic field inputs to device shape using a Bezier curve representation,
@@ 2.1.1 FBG-Based Shape Sensing with Multi-Core Fiber
Once the curvature and curvature bending angle profiles were obtained, the MATLAB script reconstructed the 3D shape of the soft robot’s centerline by applying a piecewise constant curvature (PCC) approach between each grating. The total sensing length was discretized into 25 uniform segments, each corresponding to the interval between adjacent FBGs with a fixed arc length of 10 mm. For each segment, if the curvature $\kappa$ was close to zero, it was treated as a straight section with a pure translation along the local axial direction. When curvature was present, the segment was modeled as a circular arc, where the radius was defined as $r=1/\kappa$ and the bending angle $\theta$ was computed as $\kappa\cdot ds$. A local rotation matrix was constructed for each segment to align the bending direction according to $\theta_{b}$, and the corresponding transformation was applied. These local transformations, consisting of rotation and translation, were sequentially propagated along the length of the fiber to build the global shape.
To mitigate numerical drift accumulated through successive matrix multiplications, singular value decomposition (SVD) was applied at each step to maintain orthogonality of the rotation matrices. The 3D Cartesian coordinates of the robot’s centerline were extracted from the origins of the cumulative transformation matrices, resulting in a set of 26 points representing the reconstructed backbone. Finally, the total arc length of the reconstructed shape was computed to ensure that the final shape accurately matched the known physical length of 250 mm. This reconstruction process effectively converts raw optical strain data into a precise 3D representation of the soft robot’s shape. By leveraging the PCC assumption, the algorithm achieves a balance between computational efficiency and sufficient accuracy for real-time surgical navigation. The reconstructed centerline serves as the basis for fitting a Bezier curve, providing a compact and smooth representation suitable for integration into the learning-based control framework.
@@ 3 Experimental Setup
4 EVALUATIONS AND DISCUSSION
This section presents a comprehensive evaluation of the learning based approaches used for modeling the soft suction device. We compare the performance of Neural Network (NN) and Random Forest (RF) models in predicting Bezier control points from magnetic field parameters, analyze feature importance, and evaluate shape reconstruction accuracy in the following sections.
@@ 4.1 Dataset Characteristics and Model Training
The dataset consists of 5,097 samples representing various device configurations under different magnetic field conditions. We randomly split the data into training (80%) and testing (20%) sets. The input features include $b_{x}$, $b_{y}$, magnetic field magnitude, actuation frequency, and tip distance from the coil system. For the NN model, we implemented a feedforward architecture with three hidden layers (64, 32, and 16 neurons) using the Levenberg-Marquardt backpropagation algorithm. As shown in Fig. 7, the model achieved optimal performance at approximately 30% of the training data, with minimal improvement thereafter, suggesting efficient learning from a relatively small portion of the dataset. The overall correlation coefficient across all 12 output variables was $R=0.90574$. For the RF model, we trained 12 separate regressors using 200 trees per forest with a minimum leaf size of 5.
Figure 7: Neural Network learning curve analysis showing the relationship between training data size and model performance.
@@ 4.2 Model Performance Comparison
Table 2 presents the detailed performance metrics of both models across all control points. The RF model outperforms the NN model, achieving an overall 28.37% reduction in RMSE. This improvement is particularly pronounced for the distal control points ($\mathbf{p}_{2}$ and $\mathbf{p}_{3}$), which are critical for accurate device tip positioning. The model accuracy comparison by individual control points, as shown in Figs. 8 and 9, reveals that while both models predict proximal control points ($\mathbf{p}_{0}$ and $\mathbf{p}_{1}$) with similar accuracy, the RF model outperforms the NN model for distal control points ($\mathbf{p}_{2}$ and $\mathbf{p}_{3}$). Quantitatively, the RF model reduces RMSE by 5.97% for $\mathbf{p}_{0}$, 5.47% for $\mathbf{p}_{1}$, 16.41% for $\mathbf{p}_{2}$, and an impressive 53.94% for $\mathbf{p}_{3}$ compared to the NN model. The error distribution analysis reveals that both models exhibit primarily zero-centered errors (Fig. 10), with the RF model demonstrating tighter error distributions across all control points. The NN model struggles with $\mathbf{p}_{3}$ coordinates, exhibiting error exceeding 1.5 mm for ${p}_{3Y}$, while the RF model maintains max absolute errors below 1.11 mm.
Table 2: Performance comparison between Neural Network (NN) and Random Forest (RF) models.

### v3

@@ Abstract
This paper introduces a learning-based modeling framework for a magnetically steerable soft suction device designed for endoscopic endonasal brain tumor resection. The device is miniaturized (4 mm outer diameter, 2 mm inner diameter, 40 mm length), 3D printed using biocompatible SIL 30 material, and integrates embedded Fiber Bragg Grating (FBG) sensors for real-time shape feedback. Shape reconstruction is represented using four Bezier control points, allowing for a compact and smooth representation of the device’s deformation. A data-driven model was trained on 5,097 experimental samples to learn the mapping from magnetic field parameters (magnitude: 0–14 mT, frequency: 0.2–1.0 Hz, and vertical tip distances from the surface of the electromagnet coil table: 90–100 mm) to the resulting geometric configuration of the soft robot, represented by four Bezier control points that define its 3D shape. The model was implemented and compared using both Neural Network (NN) and Random Forest (RF) architectures. The RF model outperformed the NN across all metrics, achieving a mean root mean square error of 0.087 mm in control point prediction and a mean shape reconstruction error of 0.064 mm. Feature importance analysis further revealed that magnetic field components predominantly influence distal control points, while frequency and distance affect the base configuration. Unlike prior studies that apply general machine learning methods to soft robotic data, the proposed framework introduces a new modeling paradigm that links magnetic actuation inputs directly to geometric Bezier control points, creating an interpretable and low-dimensional representation of deformation. This conceptual integration of magnetic field characterization, embedded FBG sensing, and Bezier-based learning provides a unified modeling strategy that can be extended to other magnetically actuated continuum robots. This learning-based approach effectively models the complex nonlinear behavior of hyperelastic soft robots under magnetic actuation without relying on simplified physical assumptions. By enabling sub-millimeter shape prediction accuracy and real-time inference, this work establishes an advancement toward the intelligent control of magnetically actuated soft robotic tools in minimally invasive neurosurgery.
@@ 1 Introduction
Figure 1: Magnetically steerable soft suction device for endoscopic endonasal brain tumor resection. (Top) Conceptual illustration showing how the device is inserted through the nasal cavity and sphenoid sinus using a standard suction endoscope to reach deep-seated skull-base regions. (Bottom) Cross-sectional schematic illustrating the device’s mechanical design and dimensions. The soft body (gray) encloses a suction lumen (2mm) and a channel routing the multi-core Fiber Bragg Grating (FBG) sensor at the neutral axis for accurate curvature sensing. The outer diameter is 4mm, and the active flexible length is 40mm. A hollow cylindrical permanent magnet (red/blue) at the distal tip enables magnetic steering under externally applied rotating magnetic fields. Additional geometric and material specifications are provided in Table 1.
@@
Traditional analytical models and constant-curvature kinematic formulations have been widely used to derive closed-form relationships between actuation inputs and the resulting configuration of continuum and soft robots [42, 34]. The constant-curvature assumption, as systematically formalized by Webster and Jones [42], provides computationally efficient closed-form solutions by treating each robot segment as a circular arc with uniform bending. Godage et al. [14] extended this framework through modal kinematics for multisection continuum arms, enabling coordinated control of multiple bending segments. While effective for simple bending scenarios, these models rely on simplifying assumptions such as uniform curvature, isotropic stiffness, and quasi-static actuation. More advanced physics-based frameworks have been developed to address these limitations. Cosserat rod theory, as applied to continuum robots by Rucker and Webster [34], provides geometrically exact formulations that capture all six deformation modes (bending about two axes, torsion, extension, and shear about two axes). Till et al. [38] demonstrated real-time dynamics simulation of Cosserat-based models using implicit integration schemes, achieving computational efficiency suitable for control applications. Building upon these foundations, the Geometric Variable Strain (GVS) approach [26] represents the current state of the art in soft robot modeling, generalizing constant-curvature kinematics by parameterizing the continuous strain field of Cosserat rods using customizable basis functions, including polynomials, Legendre bases, or even Bezier curves. The GVS framework provides a unified Lagrangian formulation for both static and dynamic analysis of soft-rigid hybrid systems, and viscoelastic constitutive laws can be incorporated to capture rate-dependent material behavior. Della Santina et al. [8] provided a comprehensive structured taxonomy of these modeling approaches, systematically categorizing methods by their underlying mathematical assumptions and clarifying the theoretical relationships and trade-offs between model fidelity, computational cost, and generalizability.
Despite these theoretical advances, practical implementation of physics-based models for magnetically actuated soft surgical devices presents specific challenges. First, the GVS framework and similar approaches require accurate identification of constitutive parameters including Young’s modulus, shear modulus, and damping coefficients. For novel 3D-printed silicone elastomers such as SIL 30, these parameters are difficult to characterize precisely, may vary with print orientation and cure conditions, and can exhibit batch-to-batch variability inherent in additive manufacturing. Second, magnetically actuated robots operating under rotating fields generated by multi-coil electromagnetic systems experience configuration-dependent torques where the magnetic loading changes as the device deforms, creating coupled magnetoelastic boundary conditions that must be explicitly formulated and updated during simulation. Third, the spatial variation of magnetic fields across the coil workspace introduces additional complexity for field mapping and real-time computation.
For instance, constant-curvature models assume uniform bending along discrete segments, yet magnetic actuation generates spatially distributed torques that vary with the local magnetic field orientation and gradient, resulting in non-uniform curvature profiles [20, 38]. Furthermore, soft silicone materials exhibit significant viscoelastic behavior that causes frequency-dependent deformation, and accurate knowledge of material parameters such as Young’s modulus and damping coefficients is difficult to obtain for hyperelastic materials fabricated via 3D printing [41]. The consequences of modeling errors are particularly critical for neurosurgical applications, where inaccurate shape prediction can lead to positioning errors at the device tip. Even millimeter-level deviations risk damaging delicate neurovascular structures such as the optic nerves, carotid arteries, or brainstem. Moreover, real-time closed-loop control, which is essential for safe surgical navigation in confined anatomical spaces, requires computationally efficient models that can be evaluated at high frequency.
To address these practical challenges, recent studies have increasingly adopted data-driven and learning-based modeling strategies for soft and continuum robots. These frameworks use machine learning algorithms to learn the nonlinear mapping between actuation inputs and resulting deformation directly from experimental data, bypassing the need for explicit material characterization and constitutive model calibration. As discussed by Wang et al. [41], learning-based models address key challenges in soft robot modeling, including distributed compliance, material nonlinearity, and environmental interactions. Similarly, Falotico et al. [12] demonstrated that learning controllers can adapt to dynamic conditions and fabrication variability while maintaining high accuracy and real-time computational performance.
Beyond these review articles, several primary studies have demonstrated the effectiveness of learning-based approaches for soft robot shape prediction and control. Almanzor et al. [3] achieved static shape control of soft continuum robots using deep visual inverse kinematic models, showing that neural networks can effectively learn complex mappings between task-space targets and actuator inputs without explicit kinematic derivation. George Thuruthel et al. [23] introduced visuo-dynamic self-modeling for soft robotic systems, demonstrating that learning-based approaches can capture time-variant material properties and dynamic behaviors, including viscoelasticity and hysteresis, that are inherently difficult to incorporate in analytical models. These works provide compelling evidence that data-driven methods offer practical advantages for soft robot modeling, particularly when material properties are difficult to characterize or when complex dynamic effects dominate system behavior.
It is important to note that learning-based approaches are complementary to, rather than replacements for, physics-based frameworks such as GVS. While physics-based models offer superior interpretability, generalizability to untrained conditions, and capability for design optimization, learning-based methods provide a pragmatic pathway for rapidly characterizing device behavior when comprehensive material testing and magnetic field-body coupling formulation are not feasible. Future hybrid approaches that combine the GVS structure with learning-based parameter identification or model correction represent a promising research direction [26]. Building upon these insights, this work extends the learning paradigm to magnetically actuated soft surgical instruments by introducing a data-driven framework that maps magnetic field parameters to geometric Bezier control points. Unlike prior learning-based studies that primarily address tendon-driven or pneumatic soft robots, our framework specifically targets the unique challenges of magnetic actuation, where the loading depends on the spatial magnetic field distribution and the device’s current configuration. To the best of our knowledge, this represents the first application of learning-based shape prediction specifically tailored for magnetically actuated soft surgical devices with embedded FBG sensing. This compact parametric representation captures complex, spatially coupled effects such as magnetic torque distribution, viscoelastic response, and boundary constraints by training on experimental data collected across multiple actuation frequencies and field strengths.
Table 1: Geometric and material specifications of the magnetically steerable soft suction device.
Parameter Specification
Outer diameter 4 mm
Inner suction channel 2 mm diameter
Active flexible length 40 mm
Tip magnet size OD 6.35 mm, ID 3.18 mm, thickness 3.18 mm
Tip magnet magnetization Through-thickness (axial)
Material Biocompatible silicone elastomer (SIL 30)
Embedded sensor Multi-core Fiber Bragg Grating (FBG)
FBG fiber outer diameter 0.38 mm
Sensing length used 40 mm (distal segment)
@@
Recognizing the potential of combining soft robotics and magnetic actuation, a prior study in [32] demonstrated the feasibility of integrating a magnetically steerable soft tip with a conventional rigid suction device to overcome the limitations of straight, rigid tools in endoscopic procedures. The prototype in that study featured soft segments made from Ecoflex 00-50 embedded with a permanent magnet, and it achieved controlled in-plane bending angles exceeding 90 under moderate magnetic fields (below 25 mT). The device was fabricated using mold-based casting techniques, offering both sufficient suction capability and flexibility. Validation experiments in simulated neurosurgical environments, using agar-based brain phantoms, demonstrated that the device could safely access target sites and perform precise tumor extraction, highlighting its potential to improve the safety and efficacy of minimally invasive neurosurgical interventions.
@@
Santina [8] Soft robot modeling Analytical Structured taxonomy; comprehensive framework comparison Review/taxonomy paper
Mathew [26] Hybrid soft-rigid robots Analytical Geometrically exact; customizable strain bases; unified dynamic framework Requires material parameter identification
Almanzor [3] Shape control Learning Deep visual inverse kinematics; task-space control Requires visual feedback
Thuruthel [23] Dynamic modeling Learning Visuo-dynamic self-modeling; captures time-variant properties Vision-based; system-specific
@@
Shape estimation in flexible robots can generally be achieved using either sensor-free or sensor-based methods. Sensor-free approaches rely on kinematic or magnetic models to infer shape without embedded sensors [18, 31], while sensor-based techniques use embedded elements for direct measurement [25, 39]. Our method follows the sensor-based approach, combining FBG data with a learning-based model to enable accurate, real-time shape prediction. More importantly, a learning-based modeling framework is proposed to replace the traditional mechanical models. The proposed framework directly maps magnetic field inputs to the device’s geometric configuration using Bezier control points, providing a compact and interpretable representation of the robot shape. This data-driven approach eliminates the need for explicit mechanical assumptions and instead leverages experimental data to capture the coupled nonlinear behavior between magnetic actuation and hyperelastic deformation. By combining physical sensing through embedded FBGs with data-driven modeling, the framework achieves accurate and real-time shape prediction suitable for closed-loop surgical control applications. This enables precise shape prediction, improved control, and adaptability across diverse surgical scenarios. To the best of our knowledge, this is the first magnetically actuated soft suction device equipped with a learning-based shape prediction model tailored for neurosurgical applications. The key contributions are:
@@
    Introduction of a novel learning-based modeling paradigm that connects magnetic actuation inputs directly to Bezier control points, providing a compact, interpretable, and low-dimensional representation of the device geometry suitable for real-time control,
@@
    Conceptual integration of magnetic field characterization, embedded FBG shape sensing, and Bezier-based learning into a unified modeling framework that can be generalized to other magnetically actuated soft robotic systems,
  • 4.
@@ 2.1.1 FBG-Based Shape Sensing with Multi-Core Fiber
Once the curvature and curvature bending angle profiles were obtained, the MATLAB script performed continuous shape reconstruction of the soft robot’s 3D centerline by applying piecewise constant curvature (PCC) interpolation between each grating. The total sensing length was discretized into 25 uniform segments, each corresponding to the interval between adjacent FBGs with a fixed arc length of 10 mm. For each segment, if the curvature $\kappa$ was close to zero, it was treated as a straight section with a pure translation along the local axial direction. When curvature was present, the segment was modeled as a circular arc, where the radius was defined as $r=1/\kappa$ and the bending angle $\theta$ was computed as $\kappa\cdot ds$. A local rotation matrix was constructed for each segment to align the bending direction according to $\theta_{b}$, and the corresponding transformation was applied. These local transformations, consisting of rotation and translation, were sequentially propagated along the length of the fiber to build the global shape.
To mitigate numerical drift accumulated through successive matrix multiplications, singular value decomposition (SVD) was applied at each step to maintain orthogonality of the rotation matrices. The 3D Cartesian coordinates of the robot’s centerline were extracted from the origins of the cumulative transformation matrices, resulting in a set of 26 points representing the reconstructed backbone. Finally, the total arc length of the reconstructed shape was computed to ensure that the final shape accurately matched the known physical length of 250 mm. This reconstruction process effectively converts raw optical strain data into a precise 3D representation of the soft robot’s shape. By leveraging the PCC assumption, the algorithm achieves a balance between computational efficiency and sufficient accuracy for real-time surgical navigation. The reconstructed centerline serves as the basis for fitting a Bezier curve, providing a compact and smooth representation suitable for integration into the learning-based control framework. It is important to note that this reconstruction step employs continuous shape reconstruction to obtain a smooth 3D centerline from discrete FBG curvature measurements, implemented via piecewise constant curvature (PCC) interpolation between grating nodes, which is a standard practice in FBG-based shape sensing [2, 10]. This is fundamentally different from the “traditional PCC-based modeling” discussed in the Introduction, which refers to analytical kinematic models that derive closed-form actuation-to-shape relationships under idealized constant-curvature assumptions. Such analytical models fail to capture spatially distributed magnetic torques, material viscoelasticity, and dynamic coupling effects. In contrast, our learning-based framework imposes no constant-curvature constraint on the predicted shapes. The Bezier curve fitting applied in the following section provides a smooth and compact geometric representation that abstracts the overall deformation into a small number of control points. These control points are then used as learning targets, encoding the experimentally measured deformation under each actuation condition. By training on data collected across multiple frequencies (0.2–1.0 Hz) and magnetic field strengths (0–14 mT), the model learns the actual nonlinear, frequency-dependent, and viscoelastic behavior of the soft robot, effects that are inherently captured in the experimental measurements but would be absent in analytical PCC-based kinematic models. We acknowledge that the ground truth accuracy is bounded by the FBG reconstruction fidelity; however, the learning model’s ability to capture dynamic and coupled effects from experimental data represents a fundamental advantage over traditional analytical approaches.
@@ 2.3 Soft Robot Characterization
3 Results and Discussion
This section presents a comprehensive evaluation of the learning based approaches used for modeling the soft suction device. The comparison between the Neural Network (NN) and Random Forest (RF) models was selected to represent two complementary learning paradigms for nonlinear regression. The NN provides a parametric approach suitable for learning smooth mappings between magnetic field inputs and geometric outputs, while the RF offers a non-parametric ensemble method that is more robust to noise and better suited for moderate-sized experimental datasets. Since the dataset used in this study consists of 5,097 experimentally collected samples with inherent variability and measurement noise, comparing these two models enables assessment of their relative accuracy, stability, and computational efficiency for predicting Bezier control points in real time. We compare the performance of both models in predicting Bezier control points from magnetic field parameters, analyze feature importance, and evaluate shape reconstruction accuracy in the following sections.
@@ 3.1 Dataset Characteristics and Model Training
The dataset consists of 5,097 samples representing various device configurations under different magnetic field conditions. We randomly split the data into training (80%) and testing (20%) sets. The input features used for model training include the in-plane magnetic field components $b_{x}$ and $b_{y}$, which describe the direction of the applied magnetic field, the corresponding magnetic field magnitude $|\mathbf{b}|$ representing its overall strength, the actuation frequency, and the tip distance from the coil system. The target outputs correspond to the three-dimensional coordinates of the four Bezier control points ($\mathbf{p}_{0}$$\mathbf{p}_{3}$) that parameterize the robot’s reconstructed shape from FBG measurements. These control points provide a compact geometric representation of the overall device configuration, serving as the prediction targets for both the NN and RF models. For the NN model, we implemented a feedforward architecture with three hidden layers (64, 32, and 16 neurons) using the Levenberg-Marquardt backpropagation algorithm. As shown in Fig. 7, the model achieved optimal performance at approximately 30% of the training data, with minimal improvement thereafter, suggesting efficient learning from a relatively small portion of the dataset. The overall correlation coefficient across all 12 output variables was $R=0.90574$. For the RF model, we trained 12 separate regressors using 200 trees per forest with a minimum leaf size of 5. The RF approach was selected for its robustness to measurement noise, low overfitting risk, and ability to model nonlinear input–output relationships without requiring smoothness or continuity assumptions. Each decision tree was trained on randomly sampled subsets of data and features to improve model diversity and generalization. This ensemble structure enables the RF to capture localized nonlinearities between magnetic actuation inputs and Bezier control points while maintaining computational efficiency during inference. In addition, the RF provides feature importance metrics that help identify which magnetic field components most strongly influence the robot’s deformation. Compared to the NN model, which requires careful hyperparameter tuning and larger datasets, the RF achieves high predictive accuracy with minimal tuning, making it well-suited for real-time prediction and control in surgical applications.
Figure 7: Neural network training diagnostics. (Top) Learning-curve analysis showing mean squared error versus training data size. Performance improves quickly up to roughly 30% of the data and then levels off, indicating efficient learning from a relatively small subset. (Bottom) Training and validation loss across epochs (log scale) for a feedforward network with hidden layers of 64, 32, and 16 neurons trained using Levenberg–Marquardt. Both losses drop sharply in the first few epochs and then stabilize, indicating fast convergence with consistent generalization.
@@ 3.2 Model Performance Comparison
Table 3 presents the detailed performance metrics of both models across all control points. As summarized in the table, the NN baseline achieved an overall RMSE of 0.121 mm (R2 = 0.788, max error = 1.56 mm), whereas the RF achieved 0.087 mm (R2 = 0.816, max error = 1.11 mm). Based on these absolute values, the RF model demonstrates a 28.37% reduction in RMSE compared with the NN model. This improvement is particularly pronounced for the distal control points ($\mathbf{p}_{2}$ and $\mathbf{p}_{3}$), which are critical for accurate device tip positioning. The model accuracy comparison by individual control points, as shown in Figs. 8 and 9, reveals that while both models predict proximal control points ($\mathbf{p}_{0}$ and $\mathbf{p}_{1}$) with similar accuracy, the RF model outperforms the NN model for distal control points ($\mathbf{p}_{2}$ and $\mathbf{p}_{3}$). Quantitatively, the RF model reduces RMSE by 5.97% for $\mathbf{p}_{0}$, 5.47% for $\mathbf{p}_{1}$, 16.41% for $\mathbf{p}_{2}$, and an impressive 53.94% for $\mathbf{p}_{3}$ compared to the NN model. The error distribution analysis further shows that both models exhibit primarily zero-centered errors (Fig. 10), with the RF model demonstrating tighter error distributions across all control points. The NN model struggles with $\mathbf{p}_{3}$ coordinates, exhibiting errors exceeding 1.5 mm for ${p}_{3Y}$, whereas the RF model maintains maximum absolute errors below 1.11 mm. These findings confirm that the RF model provides superior predictive accuracy, particularly in the distal region of the device where precise control is most critical.
Table 3: Performance comparison between NN and RF models.